# Knowledge Bridger: Towards Training-Free Missing Modality Completion

### v3

@@ Abstract
Previous successful approaches to missing modality completion rely on carefully designed fusion techniques and extensive pre-training on complete data, which can limit their generalizability in out-of-domain (OOD) scenarios. In this study, we pose a new challenge: can we develop a missing modality completion model that is both resource-efficient and robust to OOD generalization? To address this, we present a training-free framework for missing modality completion that leverages large multimodal model (LMM). Our approach, termed the “Knowledge Bridger”, is modality-agnostic and integrates generation and ranking of missing modalities. By defining domain-specific priors, our method automatically extracts structured information from available modalities to construct knowledge graphs. These extracted graphs connect the missing modality generation and ranking modules through the LMM, resulting in high-quality imputations of missing modalities. Experimental results across both general and medical domains show that our approach consistently outperforms competing methods, including in OOD generalization. Additionally, our knowledge-driven generation and ranking techniques demonstrate superiority over variants that directly employ LMMs for generation and ranking, offering insights that may be valuable for applications in other domains.
@@ 1 Introduction
Based on these insights, we propose a novel, training-free MMC approach, termed “Knowledge Bridger”, which autonomously mines multimodal knowledge, generates missing modalities, and ranks the best completions. This method comprises three main modules: a knowledge modeling module, a knowledge-driven modality generation module, and a ranking module.
In the knowledge modeling stage, we employ LMM to analyze available modalities and extract key elements such as objects, interactions, and attributes using the CoT approach. For specialized fields like medical imaging, we use in-context learning to integrate domain-specific knowledge, reducing reliance on extensive target domain data. Following knowledge extraction, we construct a knowledge graph to represent the relationships and attributes of available and missing modalities, providing LMM with a structured reference for generating missing content. To ensure high-quality completion, the knowledge-driven generation module utilizes this graph to guide LMM in generating specific content with precise details, such as object locations and observable attributes. Finally, the ranking module, informed by expert knowledge, assesses each generated completion by computing graph and representational similarity scores. Specifically, a graph similarity score is derived from comparing the knowledge graphs of the available and generated modalities, while a representational similarity score is calculated using models like BLIP [28] and CLIP [45]. A weighted average of these scores offers the final assessment.
@@ 3.1 Setup
Implementation Details. We employ Qwen2-VL-7B [55] as our default large multimodal model222More implementation details are presented in the appendix.. It is specifically designed to handle both visual and textual inputs, enabling it to process and interpret complex multimodal content. For image reconstruction, we apply Stable Diffusion XL (SDXL) 1.0 [43] as the restoration module for general domains. SDXL 1.0 is an advanced text-to-image diffusion model that can generate images according to a given prompt. Additionally, for the restoration of chest X-ray modality, we use Cheff [58], a cascaded chest X-ray latent diffusion model. By default, we generate 5 candidates for the missing modality during the generation process. Since the chosen LMM has already been pre-trained on the COCO dataset, to ensure a fair comparison, we employ pre-trained CLIP [45] vision and text encoders to replace the modality backbones of the comparison methods. We conduct all experiments on the PyTorch 2.4.0 [41] platform, running on Ubuntu 20.04 LTS utilizing 4 GPUs (NVIDIA GeForce RTX 4090 with 24 GB of memory). In our setting, we conduct the missing rate to simulate the missing modality scenario during training.
@@ 3.4 Visualization Analysis
To better understanding the differences between our approach and conditional generation, we present completion results produced by our method, as shown in Fig. 3. We provide visualization results across different missing modalities for three datasets555More visualization are presented in appendix.. In the general domain, our knowledge modeling module focuses on understanding the quantity of objects, their attributes, and the contextual environment. The results in Fig. 3a and Fig. 3b indicate that our method is more similar with the original missing modality compared to direct generation approaches.
In the medical domain, incorporating knowledge of different lesions enables the LMM to understand the relationships between various regions in chest X-rays and the content described by the modality. Fig. 3c and Fig. 3d show that our method provides a more reliable strategy for missing data completion than direct generation.
@@ 4.1 Missing Multimodal Learning
Missing multimodal learning enables models to train and inference effectively even when some modalities are absent. The missing multimodal learning methods can be categorized into imputation-based and non-imputation-based approaches. Imputation-based methods predict or recover missing data by learning interactions between modalities. Typical strategies include filling with special values such as zero [40] or the average value [62], which are unsuitable for high-dimensional data. Advanced methods like SMIL [36] use meta-learning to simulate missing modalities. Wang et al. [53] propose a multitask framework for missing modality learning, which learns shared features across different tasks and uses these shared features to reconstruct missing modalities. Lian et al. [30] use graph neural networks to explore relationships between modalities and recover missing ones, while Zhao et al. [65] and Pham et al. [42] leverage cycle consistency to capture cross-modal interactions for predicting missing modalities.
@@ 4.3 Conditional Generation
Conditional generation involves the creation of new data samples based on specific input data or attributes, allowing models to produce outputs tailored to given prompts or features. Early foundational work in this area includes Generative Adversarial Networks (GANs) [15] and Variational Autoencoders (VAEs) [23], with notable advancements such as conditional GANs [38] and Conditional VAEs (CVAEs) [49]. Recent progress has introduced diffusion models [50, 29, 48], which have demonstrated the ability to generate more realistic content, as exemplified by systems like DALL-E [46] and Stable Diffusion [43]. Zhang et al. introduce ControlNet [64], which significantly enhances user control over generated content by utilizing detailed, structured prompts. The key distinction between these approaches and our method lies in their respective objectives. While the former focuses on generating diverse content in response to prompts, Our method goes beyond this goal by requiring the model to understand interactions between modalities to accurately complete missing content.
@@ 6 Conclusion
Supplementary Material
G Prior Knowledge Rules and Prompts
We report the knowledge extraction rules used in our method and the prompts used in the knowledge-driven generation respectively.
G.1 Prior Knowledge
To simplify the research problem, we utilize only a set of basic instructions to direct the large language model (LMM) focus toward the desired multimodal content extraction. For the general domain, we predefine the extracted knowledge to include major objects, the quantity of each object, and their corresponding attributes and styles. Given the strong reasoning capabilities of LMM in the general domain, it is not necessary to differentiate between input modalities. However, to mitigate hallucinations generated by the LMM, we limit the number of major objects. In our experiments, extracting between 5 to 7 objects show optimal.
In the medical domain, it is necessary to pay attention to the distinctions between different modalities. This is because LMM may not inherently comprehend the knowledge required for out-of-domain (OOD) scenarios, necessitating clear identification of the modality being processed and the content to be understood. In our approach, for X-rays, we instruct the LMM to focus on anatomical structures, clinical significance, abnormal findings, and report generation.
For reports, in addition to the information extracted from X-rays, we direct the LMM to consider the locations of the anatomical structures mentioned and any additional characteristic details present in the report.
G.2 Knowledge Extraction with Chain-of-Thought
We employ an LMM with Chain-of-Thought (CoT) [60] reasoning to extract knowledge according to the aforementioned rules. This approach helps reduce the computational strain associated with long-problem reasoning and enhances the overall accuracy of problem-solving. For the general domain, our final instruction prompt is designed as follows:
For medical domain, we have:
Next, we require the LMM to integrate the CoT results according to the following instructions:
We strictly require the LMM to return structured information in the following format:
After extracting the aforementioned structured information, we employ LMM to transform these knowledge into the form of a knowledge graph. To simplify the process, we represent relationships on the graph using a triplet structure (nodes and edges):
G.3 Knowledge-driven Generation
We employ the knowledge graphs extracted by the aforementioned process as input and employ the LMM to process this information to generate meaningful descriptions of missing modalities. Various modality generators are then utilized to produce the missing information based on these descriptions. For missing image, predictions are based on observable text:
For missing text, the same method is applied to generate descriptions of the missing content, which are then refined by the LMM to produce the required missing text.
G.4 Knowledge-based Ranking Pseudo-code
The pseudo-code for the ranking process is shown in Alg. 1.
H Implementation Details
GPUs Details. We conduct all experiments on the PyTorch 2.4.0 [41] platform, running on Ubuntu 20.04 LTS utilizing 4 GPUs (NVIDIA GeForce RTX 4090 with 24 GB of memory).
Deploy Efficient Large Multimodal Model. We deploy the Qwen-VL[55] large model using vLLM [25]. vLLM is an LLM serving system that achieves (1) near-zero waste in KV cache memory and (2) flexible sharing of KV cache within and across requests to further minimize memory usage. We deploy versions with 2B777https://huggingface.co/Qwen/Qwen2-VL-2B-Instruct, 7B888https://huggingface.co/Qwen/Qwen2-VL-7B-Instruct, and 72B parameters. Specifically, due to hardware constraints, we utilize the 72B quantized version with Int-8 precision available on Hugging Face: Qwen2-VL-72B-Instruct-GPTQ-Int8999https://huggingface.co/Qwen/Qwen2-VL-72B-Instruct-GPTQ-Int8. For all versions, we maintain an 8K context window length and support a maximum of four image queries. For each query, we set the maximum number of tokens to 512 and use a temperature of 0.1.
Generators Settings. For image reconstruction, we apply Stable Diffusion XL (SDXL) 1.0 [43] as the restoration module for general domains. SDXL 1.0 is an advanced text-to-image diffusion model that can generate images according to a given prompt. Additionally, for the restoration of chest X-ray modality, we use Cheff [58], a cascaded chest X-ray latent diffusion model. By default, we generate 5 candidates for the missing modality during the generation process.
Missing Modality Simulation In our setting, we conduct the missing rate to simulate the missing modality scenario during training. Specifically, we calculate the number of missing samples in different datasets under a given missing rate and then randomly mark the text and image modalities of these samples as missing with a probability of 0.5. To ensure the reproducibility of the experiment, we perform multiple simulations using the same set of missing samples. Finally, we retrain the baseline model, which was initially trained on complete modalities, using the data with imputed missing modality and report the performance across various metrics.
I More Experimental Results
I.1 Table of Quantitative Results
We present additional quantitative analysis results, as shown in Table 5.
| Missing Rate | 0.5 | ||
| Method | F1 | AP | SS |
| Baseline (complete) | F1: 57.0 AP: 75.7 SS: - | ||
| Baseline (remove missing) | 42.1 | 63.9 | - |
| MPMM [26] (CVPR’23) | 42.8 | 64.2 | - |
| MPLMM [17] (ACL’24) | 43.3 | 65.9 | - |
| MMIN [65] (ACL’21) | 34.8 | 60.1 | 17.0 |
| DiCMoR [56] (CVPR’23) | 37.6 | 63.8 | 17.7 |
| Ours (Qwen-VL-2B) | 46.6 | 67.1 | 21.1 |
| Ours (Qwen-VL-7B) | 50.9 | 70.2 | 21.5 |
| Complete Baseline | -6.1 | -5.5 | - |
| SOTA | +7.6 | +4.3 | +3.8 |
I.2 Table of Ablation Study
We present complete results of the ablation study in Table 6.
| MM-IMDb | IU X-ray | |||||||||||||||||
| Missing Rate | 0.3 | 0.5 | 0.7 | 0.3 | 0.5 | 0.7 | ||||||||||||
| Variants | F1 | AP | SS | F1 | AP | SS | F1 | AP | SS | F1 | AP | SS | F1 | AP | SS | F1 | AP | SS |
| Baseline (Qwen-VL-7B) | 54.7 | 60.9 | 33.5 | 54.9 | 61.3 | 32.7 | 55.2 | 61.8 | 32.3 | 53.6 | 73.9 | 22.6 | 50.9 | 70.2 | 21.5 | 46.3 | 70.5 | 19.8 |
| w/o Knowledge Modeling | -1.2 | -3.3 | -7.0 | -1.5 | -4.1 | -8.8 | -1.3 | -3.6 | -8.8 | -12.1 | -21.6 | -10.7 | -13.3 | 26.8 | -11.6 | -17.5 | -29.2 | -13.7 |
| + Random Ranking | -1.7 | -3.5 | -7.2 | -1.6 | -4.3 | -9.0 | -1.6 | -4.1 | -9.9 | -13.9 | -26.8 | -11.4 | -14.7 | -28.1 | -12.4 | -19.3 | -31.8 | -15.0 |
| Random Ranking | -0.5 | -2.6 | -0.6 | -0.4 | -2.8 | -0.7 | -0.5 | -2.7 | -0.6 | -2.9 | -6.3 | -3.8 | -3.1 | -6.9 | -4.3 | -3.8 | -7.1 | -4.7 |
| w/o Knowledge Ranking | -0.1 | -0.4 | -0.2 | -0.2 | -1.9 | -0.4 | -0.2 | -0.8 | -0.1 | -1.3 | -3.3 | -1.5 | -2.1 | -5.4 | -3.6 | -1.9 | -5.7 | -2.1 |
| w/o Semantic Ranking | -0.3 | -1.4 | -0.2 | -1.4 | -0.2 | -0.1 | -0.2 | -1.0 | -0.3 | -0.9 | -0.4 | -0.7 | -1.3 | -2.3 | -1.1 | -2.4 | -3.3 | -1.6 |
I.3 More Modalities Results
The proposed method primarily focuses on image and text modalities to facilitate the evaluation of the proposed method, as these modalities are well-supported by the community and computationally efficient when using large vision-language models. However, our approach imposes no constraints on the modality encoders, allowing it to be easily generalized to other modalities. To validate this, we conducted experiment on a multimodal sarcasm detection dataset [11], which includes audio, vision, and text modalities. Using Unified-IO [10] as the backbone and keeping other configurations unchanged, our method demonstrated effectiveness even when extended to these modalities, as shown in Table 7.
| Missing Rate | 0.3 | 0.5 | ||||
| Method | F1 | AP | SS | F1 | AP | SS |
| Baseline (complete) | F1: 62.4 mAP: 64.7 SS: - | |||||
| Baseline (remove missing) | 57.1 | 59.3 | - | 53.8 | 54.6 | - |
| DiCMoR (CVPR’23) | 55.5 | 57.6 | 11.7 | 51.3 | 52.9 | 10.2 |
| Ours (Unified-IO 7B) | 58.3 | 60.1 | 13.3 | 54.7 | 55.8 | 11.4 |
J Visualization Analysis
J.1 Completion Results
We present additional completion results, as shown in Fig. 4, 5, and 6. In the general domain, our knowledge modeling module emphasizes understanding the quantity of objects, their attributes, and the contextual environment. The results in Fig. 4 indicate that our method more closely resembles the original missing modality compared to direct generation approaches. In the medical domain, incorporating knowledge of different lesions allows the LMM to comprehend the relationships between various regions in chest X-rays and the content described by the modality. Figs. 5 and 6 demonstrate that our method offers a more reliable strategy for missing data completion than direct generation.
J.2 Intermediate Results
J.3 Knowledge-based Ranking Results
We present partial results of knowledge-based ranking as shown in Fig. 7. Here, “available [modality]” indicates that the modality is visible, and “QS” represents the quality score of the completed missing modality combined with the observed modality. The results demonstrate that our proposed knowledge-based ranking module can effectively select relatively reasonable generated outcomes.
K Hallucinations
The hallucinations of LMM [13] refer to instances when an AI model generates content that is factually incorrect, misleading, or unsubstantiated. In our approach, hallucinations primarily stem from the limitations of the training data. We advocate for a training-free method to achieve MMC, which has the advantage of being easily deployable across various domains with minimal input of relevant domain knowledge. However, the drawback is that the lack of task-specific training can result in the model having less “common sense” compared to models trained on extensive datasets. As illustrated in Fig. 10, the absence of common sense in our approach may lead to results that deviate from expected cognitive outcomes. In recent years, RAG (Retrieval-Augmented Generation) [27, 18] has been regarded as an effective technique for mitigating hallucinations in large models. This technique provides the model with external truths, thereby reducing hallucinations during the reasoning process. In future work, we plan to incorporate RAG to enhance the robustness of our approach.
L Limitations
L.1 More Modalities
Our method focuses exclusively on image and text modalities, leaving its performance on other modalities, such as speech and depth, yet to be explored. The approach emphasizes the automatic extraction of inter-modal knowledge and the completion of missing modalities through domain knowledge. However, this focus on a limited set of modalities limits its generalizability and adaptability in real-world applications where multi-modal data often involves various types of sensory inputs. Thus, in the future, adaptation to other modalities is possible by defining a more comprehensive modality knowledge and expanding the learning framework to accommodate these new modalities. Some promising works [14, 35] show that one modality, such as image or text, can be connected to any other modality, paving the way for more inclusive and versatile multi-modal systems that handle diverse data types with high efficacy.
L.2 More Tasks and Metrics
Additionally, we observe that while our method enhances classification performance under a high missing rate (e.g., 0.7), it paradoxically results in a decrease in the similarity scores of the completed modalities. This suggests that although the model performs well in reconstructing missing data for classification tasks, the semantic alignment and quality of the generated modalities may still require significant refinement. Addressing these limitations presents an opportunity to improve the balance between classification accuracy and modality similarity. Therefore, there remains substantial potential for further exploration to develop more robust generation and ranking strategies in the future. These improvements could include incorporating advanced similarity-preserving techniques and exploring diverse evaluation metrics to assess the completeness and coherence of generated data across different tasks.

### v5

@@ Abstract
Previous successful approaches to missing modality completion rely on carefully designed fusion techniques and extensive pre-training on complete data, which can limit their generalizability in out-of-domain (OOD) scenarios. In this study, we pose a new challenge: can we develop a missing modality completion model that is both resource-efficient and robust to OOD generalization? To address this, we present a training-free framework for missing modality completion that leverages large multimodal model (LMM). Our approach, termed the “Knowledge Bridger”, is modality-agnostic and integrates generation and ranking of missing modalities. By defining domain-specific priors, our method automatically extracts structured information from available modalities to construct knowledge graphs. These extracted graphs connect the missing modality generation and ranking modules through the LMM, resulting in high-quality imputations of missing modalities. Experimental results across both general and medical domains show that our approach consistently outperforms competing methods, including in OOD generalization. Additionally, our knowledge-driven generation and ranking techniques demonstrate superiority over variants that directly employ LMMs for generation and ranking, offering insights that may be valuable for applications in other domains. Our code is accessible at: https://github.com/Guanzhou-Ke/Knowledge-Bridger.
@@ 1 Introduction
Based on these insights, we propose a novel, training-free MMC approach, termed “Knowledge Bridger”, which autonomously mines multimodal knowledge, generates missing modalities, and ranks the best completions. This method comprises three main modules: a knowledge modeling module, a knowledge-driven modality generation module, and a ranking module. In the knowledge modeling stage, we employ LMM to analyze available modalities and extract key elements such as objects, interactions, and attributes using the CoT approach. For specialized fields like medical imaging, we use in-context learning to integrate domain-specific knowledge, reducing reliance on extensive target domain data. Following knowledge extraction, we construct a knowledge graph to represent the relationships and attributes of available and missing modalities, providing LMM with a structured reference for generating missing content. To ensure high-quality completion, the knowledge-driven generation module utilizes this graph to guide LMM in generating specific content with precise details, such as object locations and observable attributes. Finally, the ranking module, informed by expert knowledge, assesses each generated completion by computing graph and representational similarity scores. Specifically, a graph similarity score is derived from comparing the knowledge graphs of the available and generated modalities, while a representational similarity score is calculated using models like BLIP [22] and CLIP [38]. A weighted average of these scores offers the final assessment.
@@ 3.1 Setup
Implementation Details. We employ Qwen2-VL-7B [48] as our default large multimodal model222More implementation details are presented in the appendix.. It is specifically designed to handle both visual and textual inputs, enabling it to process and interpret complex multimodal content. For image reconstruction, we apply Stable Diffusion XL (SDXL) 1.0 [36] as the restoration module for general domains. SDXL 1.0 is an advanced text-to-image diffusion model that can generate images according to a given prompt. Additionally, for the restoration of chest X-ray modality, we use Cheff [51], a cascaded chest X-ray latent diffusion model. By default, we generate 5 candidates for the missing modality during the generation process. Since the chosen LMM has already been pre-trained on the COCO dataset, to ensure a fair comparison, we employ pre-trained CLIP [38] vision and text encoders to replace the modality backbones of the comparison methods. We conduct all experiments on the PyTorch 2.4.0 platform, running on Ubuntu 20.04 LTS utilizing 4 GPUs (NVIDIA GeForce RTX 4090 with 24 GB of memory). In our setting, we conduct the missing rate to simulate the missing modality scenario during training.
@@ 3.4 Visualization Analysis
To better understanding the differences between our approach and conditional generation, we present completion results produced by our method, as shown in Fig. 3. We provide visualization results across different missing modalities for three datasets555More visualization are presented in appendix.. In the general domain, our knowledge modeling module focuses on understanding the quantity of objects, their attributes, and the contextual environment. The results in Fig. 3a and Fig. 3b indicate that our method is more similar with the original missing modality compared to direct generation approaches. In the medical domain, incorporating knowledge of different lesions enables the LMM to understand the relationships between various regions in chest X-rays and the content described by the modality. Fig. 3c and Fig. 3d show that our method provides a more reliable strategy for missing data completion than direct generation.
@@ 4.1 Missing Multimodal Learning
Missing multimodal learning enables models to train and inference effectively even when some modalities are absent. The missing multimodal learning methods can be categorized into imputation-based and non-imputation-based approaches. Imputation-based methods predict or recover missing data by learning interactions between modalities. Typical strategies include filling with special values such as zero or the average value, which are unsuitable for high-dimensional data [34] . Advanced methods like SMIL [30] use meta-learning to simulate missing modalities. Wang et al. [46] propose a multitask framework for missing modality learning, which learns shared features across different tasks and uses these shared features to reconstruct missing modalities. Lian et al. [24] use graph neural networks to explore relationships between modalities and recover missing ones, while Zhao et al. [57] and Pham et al. [35] leverage cycle consistency to capture cross-modal interactions for predicting missing modalities.
@@ 4.3 Conditional Generation
Conditional generation creates new data samples based on specific inputs, enabling models to generate tailored outputs. Early foundational work in this area includes Generative Adversarial Networks (GANs) [11] and Variational Autoencoders (VAEs) [19], with notable advancements such as conditional GANs [32] and Conditional VAEs (CVAEs) [42]. Recent progress has introduced diffusion models [43, 23, 41], which have demonstrated the ability to generate more realistic content, as exemplified by systems like DALL-E [39] and Stable Diffusion [36]. ControlNet [56] further improves user control. The key distinction between these approaches and our method lies in their respective objectives. Unlike these approaches, which focus on diverse content generation, our method emphasizes understanding modality interactions to accurately complete missing content.